the following system of equations is a dependent system.
4x + 5y = 10
8x + 10y = 20
what is the correct way to write the solution?
a) 0
b) all numbers
c) 4x + 5y = 10
d) (0,4)

Answer by stanbon(75887)

the following system of equations is a dependent system.
4x + 5y = 10
8x + 10y = 20
what is the correct way to write the solution?
a) 0
b) all numbers
c) 4x + 5y = 10
d) (0,4)
==================
Ans: c
Cheers,
Stan H.